OpenStudy (anonymous):

Find the discriminant of the following equation. 4x2 + 16x + 16 = 0

OpenStudy (amistre64):

the generic form of a quadratic is: ax^2 +bx + c where the discriminant is: b^2 -4ac

OpenStudy (amistre64):

i cant recall if you gotta factor out that 4 first or not

OpenStudy (anonymous):

0

OpenStudy (amistre64):

the wolf says it 0 so lets see :) 16^2 -4(16)(4) = 4^2-4(4) (4^2)^2 -4(4)(4)(4)-4^2 +4(4) = 0 4^4 - 4^4 = 0 ; guess it doesnt matter if you factor or not :)

OpenStudy (amistre64):

factored might make the numbers smaller to manage better tho

OpenStudy (amistre64):

4^2 - 4(4) = 0 ; so discriminant is 0 ;)

OpenStudy (anonymous):

16^2-4*4*16=0